OpenStudy (bellaisinhere):

Write an equation in point-slope form for the line through the given point with the given slope. (–4, –2); m = -1/2 a. y+2=-1/2(x-4) b. y+2=(-1/2)x-4 c. y+2=-1/2(x+4) d. y-2=-1/2(x-4)

OpenStudy (tkhunny):

Can you write a generic point slope form? If it passes through the point (a,b) and has slope "m", what might you create?

OpenStudy (bellaisinhere):

I really don't know

OpenStudy (tkhunny):

Okay, but you have to do better than that. You WILL BE required to know this on examination. Point Slope: (y-b) = m(x-a) If we are given the point (a,b) and the slope m. Can you substitute the three values from your problem and see what we get?

OpenStudy (bellaisinhere):

2, 1/2, 4?

OpenStudy (tkhunny):

a = -4 b = 2 n = -1/2

OpenStudy (bellaisinhere):

Thank you so much! @tkhunny

OpenStudy (tkhunny):

NP.